OpenStudy (anonymous):

I don't get how to do this one question. :/ A city has a population of 23,000 in 2005, 27,600 in 2006, 33,120 in 2007, and 39,744 in 2008. If this pattern continues, what will be the population in 2012? Write the iterative rule for the arithmetic or geometric sequence. Answer the question and show all work.

OpenStudy (bibby):

If it were an arithmetic sequence, there'd be a common difference between each year If it were a geometric sequence, each successive term would be multiplied by a common ratio the sequence is 23,000, 27,600, 33,120, 39,744 the first step is finding the ratio (or the difference if it's an arithmetic sequence), then you can either use a formula or count your way to 2012

OpenStudy (jordan123321):

bibby is right

OpenStudy (bibby):

I didn't say anything rofl

OpenStudy (anonymous):

If it's a geometric sequence, the ratio would be 12 correct? Because 276000/23000 = 12.

OpenStudy (bibby):

if it's a geometric sequence, the ratio would apply to the next 2 terms

OpenStudy (bibby):

so test that out

OpenStudy (anonymous):

So it's arithmetic. 39744/33120 =1.2

OpenStudy (bibby):

not exactly. the common ratio is 1.2, you had an extra 0 in the first number which put your ratio off by a factor of 10 :p

OpenStudy (anonymous):

Oh, wow. My bad. So it is geometric.

OpenStudy (bibby):

it is, it is. given a geometric sequence, we can calculate the n-th term using \(a_n=a_1r^{n-1}\)

OpenStudy (anonymous):

Ok so a1 would be the first term right? N is the number of the term, and r would be the ratio?

OpenStudy (bibby):

yep yep yep

OpenStudy (anonymous):

an = 23,000(1.2)^12-1

OpenStudy (anonymous):

Since it's the 12th in the sequence?

OpenStudy (bibby):

how'd you get 12? our thingie starts at 2005

OpenStudy (anonymous):

So it would be 5-1? I was thinking backwards ;-;

OpenStudy (bibby):

so if 2005 is the first, we'll do this

OpenStudy (bibby):

\(\large 1_{2005},2_{2006},3_{2007},4_{2008},5_{2009},6_{2010},7_{2011},8_{2012}\) alternative, I guess we could do 2012-2005+1, but I can't logically explain why that's the case, I just know it, so I'll avoid using it

OpenStudy (bibby):

I mean I guess the + 1 is because we're including 2005. I suck at verbalizing so just use my first statement and count the years

OpenStudy (anonymous):

So since there are eight years, it would be... an = 23,000(1.2)^8-1

OpenStudy (bibby):

a_8 but yeah

OpenStudy (anonymous):

What does a_8 mean?

OpenStudy (bibby):

the underscore is LaTeX for a subscript. It means \(a_8\)

OpenStudy (anonymous):

Would I still do 8-1? Orrr would it be something else?

OpenStudy (bibby):

yeah, yeah, sorry for the confusion recall \(\huge a_n=a_1*r^{n-1}\)

OpenStudy (anonymous):

Ok. an = 23,000(1.2)^8-1 an = 27600^8-1 an = It gave me some wack decimal. Like, 2.08827...(x10)35 ;-;

OpenStudy (bibby):

shh, no tears PEDMAS/PEMDAS first, let's do the exponent, 1.2^7=?

OpenStudy (anonymous):

3.6

OpenStudy (bibby):

closer to 3.58318. it's better to not round until you finish your calculations cause it's more accurate 23,000*3.583

OpenStudy (anonymous):

Oh alright. I just rounded b/c it's usually what we're told to do. It would be 82409

OpenStudy (bibby):

we can plug it into wolfram to check http://www.wolframalpha.com/input/?i=23%2C000%281.2%29%5E7

OpenStudy (bibby):

kinda close

OpenStudy (anonymous):

And this would be the population for 2012?

OpenStudy (bibby):

it would. let's check our work another way

OpenStudy (bibby):

\(\begin{array}{|l|c|r|c|l|} \hline 2005&23,000\\ \hline 2006&27,600\\ \hline 2007&33,120\\ \hline 2008&39,744\\ \hline 2009&?\\ \hline 2010&?\\ \hline 2011&?\\ \hline 2012&?\\ \hline \end{array} \) start at 39,744 and multiply by 1.2 to get 2009 again for 2010 again for 2011 again for 2012

OpenStudy (anonymous):

47692 57231 68677 82413

OpenStudy (bibby):

yeaaaaa boiii

OpenStudy (anonymous):

Tyyy so much. And ty for sticking with it :))

OpenStudy (bibby):

it's what I do. thanks for being a good student and actually putting forth the effort. it makes things all the more worthwhile feel free to tag me in the future @bibby

OpenStudy (anonymous):

Will do! Thanks again!